# Semi-Random Graphs, Robust Asymmetry, and Reconstruction

### v1

@@ 3.1 Asymmetry
Asymmetry properties are commonly leveraged in probabilistic reconstruction arguments. Namely, such arguments often invoke the asymmetry of all sufficiently large (induced) subgraphs of the underlying random graph. The dependence of this technique upon all subgraphs enjoying this property can occasionally be avoided by instead assuming the uniqueness of certain subgraphs. Farhadian (2017) exploited this approach, demonstrating that all graphs containing a unique asymmetric subgraph on n2n-2 vertices are reconstructible. As our first lemma, we present a mild generalization of this result, which admits a constructive proof.
@@ 3.2 Robust Asymmetry
Let us begin by articulating subgraph uniqueness properties which will play a central role in our forthcoming results. Recall that for a graph $G=(V,E)$ and subset $S\subset V^{(2)}$, $B_{S}(G)$ denotes the set of graphs generated by adding/removing any number of vertex pairs from $S$ in $G$,
@@
Informally, $E_{\delta}(G)$ records the condition that all subgraphs of $G$ on $n$ vertices are asymmetric and mutually non-isomorphic, while $E_{\delta}(G,S)$ records whether this is true even when one is permitted to add/remove any vertex pairs in $S$ from $G$. We are now equipped to state our primary lemma.
@@
$\frac{\log(n)}{n}\leq\beta(n)=o(1),\qquad\quad p(n)\in[\alpha\beta(n),1-\alpha\beta(n)]^{\binom{n}{2}}$
then for sufficiently large $n$, $G\sim\mathscr{G}(n+\delta,p)$, and independent $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon}$, $E_{\delta}(G,S)$ holds with failure probability at most $\exp(-\rho\beta(n)n)$.
@@
Here and in the remaining results, it makes no difference whether we draw $\varepsilon=\varepsilon(n)$ or $\varepsilon=\varepsilon(n+\delta)$ pairs. For simplicity in the proofs, however, we take $\varepsilon=\varepsilon(n)$. A crucial difference between the cases where $\beta=o(1)$ and $\beta=\Theta(1)$ is the adaptivity of our constants. In the former case, we can choose $\alpha$ to scale the convergence rate by any $\rho$ of our choosing. This property will be useful when we examine downstream implications.
After some simple observations on the front end, our proof of Lemma 3.4 closely follows that of Bollobás (1990). Roughly, we upper bounds the probability of any single isomorphism $\varphi$ by the probability that its orbits on $V^{(2)}$ consist entirely of edges or entirely of nonedges. We find that we can fit in additional cross-terms that account for the edits contributed by $S$. Here and in the sequel, proofs for $\beta\in\Theta(1)$ are entirely similar to those for $\beta\in o(1)$; we leave such proofs to Appendix B.
@@
$\binom{n+\delta}{\delta}\leq 2n^{\delta}.$
Fix one such $W$. Let $S_{n}^{(m)}$ denote the set of injections $W\to V$ that move exactly $m$ vertices of $W$. We have
$|S_{n}^{(m)}|\leq\binom{n}{m}(m+\delta)_{m}=n_{m}\binom{m+\delta}{\delta}\leq 2n^{m}m^{\delta}.$
For a map $\varphi\in S_{n}^{(m)}$, let $\mathscr{G}\times V^{(2)}(\varphi)$ denote the set of $(G,S)$ pairs for which $\varphi$ induces an isomorphism $G^{\prime}[W]\to G^{\prime}[\varphi(W)]$ for some $G^{\prime}\in B_{S}(G)$. We let $\operatorname*{\mathbb{P}}[\varphi]=\operatorname*{\mathbb{P}}[\mathscr{G}\times V^{(2)}(\varphi)]$. Note that $\varphi$ defines a partition of $W^{(2)}\cup\varphi(W)^{(2)}$ into orbits. Let $M_{i}$ denote the number of orbits of order $i$, and $N=\binom{n}{2}$, so that
@@
$\leq 4n^{\delta+1}\exp\left[-\frac{c_{1}c_{2}\alpha}{2}\beta(n)n\right]$ (1)
@@
We bound terms (1)-(3). Take any $\rho^{\prime}>\rho$. Choose $\alpha>\left(\frac{2}{c_{1}c_{2}}\right)(\rho^{\prime}+\delta+1)$. For sufficiently large $n$,
@@
We appeal to Lemma 3.4 in a black-box manner. Let $\rho^{\prime}>\rho$, $\varepsilon(n)\leq c\binom{n}{2}$, and abbreviate $N=\binom{n}{2}$ for concision. Then set
@@
with $\gamma=\frac{1+1/c}{2}$. It is not difficult to see that $\varepsilon^{\prime}(n)=O(N)=O(n^{2})$, as $\gamma\varepsilon(n)/N\leq\gamma\cdot c=\frac{c+1}{2}<1$. Now consider the random variable $Z_{n}=|S|$ for $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon^{\prime}(n)}$. Equivalently, $Z_{n}=\sum_{i\in[N]}X_{i}$ where $X_{i}=\mathbb{I}[i\in S]$. It suffices to demonstrate that $Z_{n}\geq\varepsilon(n)$ with probability $\Omega(1)$, from which one can invoke Lemma 3.4 with $\varepsilon^{\prime}(n)$ to conclude the argument by conditioning and bounding the failure probability by $O(\exp(-\rho^{\prime}\beta(n)n))$.
@@
as desired. The penultimate line uses that $\operatorname*{\mathbb{E}}(Z_{n})\geq 1$, and the previous line uses $\operatorname*{Var}(Z_{n})\leq\operatorname*{\mathbb{E}}(Z_{n})$. ∎
@@ 3.3 Planted Structure
Note that in the proof of Corollary 3.8, we make use the adaptive constants in Theorem 3.6. We can again use this adaptivity to recover variations on existing results in the literature on property resilience (Sudakov and Vu, 2008; Frieze and Karoński, 2015; Bollobás, 2011).
@@ 3.4 Reconstruction Consequences
Through Lemma 3.1, a number of reconstruction results follow immediately from Theorems 1.1, 1.2 and 1.3. Using techniques from Bollobás (1990), one can even show that semi-random graphs have reconstruction number three with failure probability $\exp(-\Omega(\beta(n)n))$. Perhaps more interesting is the following characterization of nonreconstructible graphs, which follows from Corollaries 3.10 and 3.11.
@@ 4 Outlook
As noted in the technical overview, when $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon}$, any reconstructibility result holding with probability 1o(1)1-o(1) and $\varepsilon=\Omega(n^{2}\log(n))$ is equivalent to the deterministic conjecture for large $n$. This leaves a logarithmic gap unknown. Similarly, when $S\sim\mathcal{U}(V^{(2,\varepsilon)})$, it remains unclear whether one can handle any $\varepsilon=n^{2}-o(n^{2})$. Such questions outline what is possibly most natural line for further inquiry.
As suggested in Section 1.1, one can view the semi-random extension of the smoothed model as, instead of uniformly widening the hypercube bound on $\textbf{{p}}\in[0,1]^{N}$, allowing a random (or adversarially chosen) subset $S$ of the entries in this tuple to be arbitrary. One might consider a relaxed setting, where instead of letting $\{p_{i}\}_{i\in S}$ be arbitrary, widening the bound on these entries beyond the $\Theta(\log(n)/n)$ barrier. One may hope to successfully do this on $n^{2}-o(n^{2})$ of the entries.
Beyond the semi-random model, there exist several natural notions of approximate graph reconstruction. For instance, take $\operatorname{d\mathfrak{R}}(G)$ to be the minimum number of edge additions/removals necessary to make $G$ reconstructible. One can always isolate a vertex to make $G$ reconstructible, giving $\operatorname{d\mathfrak{R}}(G)\leq n-1$. To the best of our knowledge, however, no sublinear bounds have been established for $\operatorname{d\mathfrak{R}}$. Interestingly, by strategically changing the degree profile and using Myrvold et al. (1987, Theorem 12), one can show that only three edits suffice to make any graph on at least four vertices edge-reconstructible—reconstructible from its multiset of subgraphs with a single edge removed (Bondy, 1991). We leave examination of these bounds to future work.
Appendix A Probability Tools
@@ A.1 Negative Association
We closely follow a standard proof for balls and bins. For $e\in V^{(2)}$ and $j\in[\varepsilon]$, let $X_{e,j}$ be a random variable taking the value 1 if $S_{j}=e$ and 0 otherwise. By the Zero-One Lemma (Dubhashi and Ranjan, 1998), $\{X_{e,j}\}_{e}$ is NA. Since unions of mutually independent NA families are NA, $\{X_{e,j}\}_{e,j}$ is NA (Joag-Dev and Proschan, 1983). But note that $\{\mathbb{I}[e\in S]\}_{e\in V^{(2)}}$ are all monotone increasing in $\{X_{e,j}\}_{e,j}$ and defined on disjoint subsets of $\{X_{e,j}\}_{e,j}$, hence NA. Using the marginal probability bound for NA families,
@@
For $e\in V^{(2)}$, also define $Y_{e}=\mathbb{I}[e\in S]$. Note that $\{X_{j}\}_{j}$ are mutually independent, and that $\{Y_{e}\}_{e\in A\cup B}$ is NA by Lemma A.3. Since these collections are independent, $\mathcal{A}=\{X_{j}\}_{j}\cup\{Y_{e}\}_{e\in A\cup B}$ is NA (Joag-Dev and Proschan, 1983). Now observe that
@@ Appendix B Omitted Proofs
$\operatorname*{\mathbb{P}}[\varphi]$ $\leq\exp\left[-\frac{c_{1}}{4}\beta(n)\sum_{i=2}^{N}iM_{i}\right]\leq\begin{cases}\exp\left[-\frac{c_{1}}{2}(n-1)\right]&m=1\\ \exp\left[-\frac{c_{1}}{4}m\left(n-\frac{m}{2}-1\right)\right]&m\geq 2.\end{cases}$
@@
$\leq 4n^{\delta+1}\exp\left[-\frac{c_{1}c_{2}}{2}n\right]$ (4)
@@
$\leq\sum_{m=\lfloor n^{1/2}\rfloor}^{n}\exp\left[m\left(-\frac{\rho^{\prime\prime}}{2}n\right)\right]\leq\exp(\log(n)-\rho^{\prime\prime}(n))\leq\exp(-\rho^{\prime}n).$

### v2

@@ 3.1 Asymmetry
Asymmetry properties are commonly leveraged in probabilistic reconstruction arguments. Namely, such arguments often invoke the asymmetry of all sufficiently large (induced) subgraphs of the underlying random graph. The dependence of this technique upon all subgraphs enjoying this property can occasionally be avoided by instead assuming the uniqueness of certain subgraphs. farhadian2017simple exploited this approach, demonstrating that all graphs containing a unique and asymmetric subgraph on n2n-2 vertices are reconstructible. As our first lemma, we present a mild generalization of this result, which admits a constructive proof.
@@
Proof.
Let $G-\{x,y\}$ denote a subgraph of $G$ that is both unique and asymmetric. Let $\sigma:G\xrightarrow{\sim}H$. By Fact 2.3, there is a unique and asymmetric subgraph $H-\{x^{\prime},y^{\prime}\}$ of $H$ such that $H-\{x^{\prime},y^{\prime}\}=H[\sigma(V(G)\setminus\{x,y\})]$. Hence, $\{\sigma(x),\sigma(y)\}=\{x^{\prime},y^{\prime}\}$.
Let $\varphi_{x}:G-x\xrightarrow{\cong}H-\sigma(x)$ and $\varphi_{y}:G=y\xrightarrow{\cong}H-\sigma(y)$. By uniqueness of $H-\{x^{\prime},y^{\prime}\}$ in $H$, $\varphi_{x}(V(G)\setminus\{x,y\})=\varphi_{y}(V(G)\setminus\{x,y\})=V(H)\setminus\{x^{\prime},y^{\prime}\}$. So $\varphi_{x}(y)=\sigma(y)$ and $\varphi_{y}(x)=\sigma(x)$. Finally, note that condition (iii) of Lemma 3.1 holds by the assumption that $G-\{x,y\}$ is asymmetric. ∎
@@ 3.2 Robust Asymmetry
Let us begin by articulating subgraph uniqueness properties that will play a central role in our forthcoming results. Recall that for a graph $G=(V,E)$ and subset $S\subset V^{(2)}$, $B_{S}(G)$ denotes the set of graphs generated by adding/removing any number of vertex pairs from $S$ in $G$,
@@
Informally, $E_{\delta}(G)$ records the condition that all subgraphs of $G$ on $n$ vertices are asymmetric and mutually non-isomorphic, while $E_{\delta}(G,S)$ records whether this is true even when one is permitted to add/remove any vertex pairs in $S$ from $G$. Note that the negation of $E_{\delta}(G,S)$ is monotonic in $S$. We are now equipped to state our primary lemma.
@@
$\frac{\log(n)}{n}\leq\beta(n)=o(1),\qquad\quad\textbf{{p}}(n)\in[\alpha\beta(n),1-\alpha\beta(n)]^{\binom{n}{2}}$
then for sufficiently large $n$, $G\sim\mathscr{G}(n+\delta,\textbf{{p}})$, and independent $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon}$, $E_{\delta}(G,S)$ holds with failure probability at most $\exp(-\rho\beta(n)n)$.
@@
Here and in the remaining results, it makes no difference whether we draw $\varepsilon=\varepsilon(n)$ or $\varepsilon=\varepsilon(n+\delta)$ pairs. For simplicity in the proofs, however, we take $\varepsilon=\varepsilon(n)$ and require that its image be contained in $\mathbb{N}$. A crucial difference between the cases where $\beta=o(1)$ and $\beta=\Theta(1)$ is the adaptivity of our constants. In the former case, we can choose $\alpha$ to scale the convergence rate by any $\rho$ of our choosing. This property will be useful when we examine downstream implications.
After some simple observations on the front end, our proof of Lemma 3.4 closely follows that of bollobas1990almost. Roughly, we upper bounds the probability of any single isomorphism $\varphi$ by the probability that its orbits on $V^{(2)}$ consist entirely of edges or entirely of nonedges. We find that we can fit in additional cross-terms that account for the edits contributed by $S$. Here and in the sequel, proofs for $\beta\in\Theta(1)$ are entirely similar to those for $\beta\in o(1)$; we leave such proofs to Appendix B.
@@
$\binom{n+\delta}{\delta}\leq 2n^{\delta},$
for $n\geq 2$. Fix one such $W$. Let $S_{n}^{(m)}$ denote the set of injections $W\to V$ that move exactly $m$ vertices of $W$. We have
$|S_{n}^{(m)}|\leq\binom{n}{m}(m+\delta)_{m}=n_{m}\binom{m+\delta}{\delta}\leq 2n^{m}m^{\delta},$
when $m\geq 2$, and a bound of $2\delta n$ when $m=1$. For a map $\varphi\in S_{n}^{(m)}$, let $\mathscr{G}\times V^{(2)}(\varphi)$ denote the set of $(G,S)$ pairs for which $\varphi$ induces an isomorphism $G^{\prime}[W]\to G^{\prime}[\varphi(W)]$ for some $G^{\prime}\in B_{S}(G)$. We let $\operatorname*{\mathbb{P}}[\varphi]=\operatorname*{\mathbb{P}}[\mathscr{G}\times V^{(2)}(\varphi)]$. Note that $\varphi$ defines a partition of $W^{(2)}\cup\varphi(W)^{(2)}$ into orbits. Let $M_{i}$ denote the number of orbits of order $i$, and $N=\binom{n}{2}$, so that
@@
$\leq 4\delta n^{\delta+1}\exp\left[-\frac{c_{1}c_{2}\alpha}{2}\beta(n)n\right]$ (1)
@@
We bound terms (1)-(3) by taking $n$ sufficiently large (independent of $m$). Let $\rho^{\prime}>\rho$. Choose $\alpha>\left(\frac{2}{c_{1}c_{2}}\right)(\rho^{\prime}+\delta+1)$.
@@
Let $\rho^{\prime}>\rho$, $\varepsilon(n)\leq c\binom{n}{2}$, and abbreviate $N=\binom{n}{2}$ for concision. Then set
@@
with $\gamma=\frac{1+1/c}{2}$. It is not difficult to see that $\varepsilon^{\prime}(n)=O(N)=O(n^{2})$, as $\gamma\varepsilon(n)/N\leq\gamma\cdot c=\frac{c+1}{2}<1$. Now consider the random variable $Z_{n}=|S|$ for $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon^{\prime}(n)}$. Equivalently, $Z_{n}=\sum_{1}^{N}X_{i}$ where $X_{i}=\mathbb{I}[i\in S]$. We start by showing $Z_{n}\geq\varepsilon(n)$ with probability $\Omega(1)$.
@@
as desired. The penultimate line uses that $\operatorname*{\mathbb{E}}(Z_{n})\geq 1$, and the previous line uses $\operatorname*{Var}(Z_{n})\leq\operatorname*{\mathbb{E}}(Z_{n})$. By Lemma 3.4 with $\varepsilon^{\prime}(n)$, we now have
[…]
and conclude by conditioning. To be explicit, note that for fixed $G$ and $1\leq k\leq\varepsilon^{\prime}-1$,
[…]
$=\sum_{|S|=k}\operatorname*{\mathbb{P}}_{S\sim V^{(2,k)}}[S]\cdot\mathbb{I}[E_{\delta}(G,S)^{c}]$
$\leq\sum_{|S|=k}\operatorname*{\mathbb{P}}_{S\sim V^{(2,k)}}[S]\cdot\operatorname*{\mathbb{P}}_{e\sim V\setminus S}[E_{\delta}(G,S\cup\{e\})^{c}]$
$=\operatorname*{\mathbb{P}}_{\begin{subarray}{c}S\sim V^{(2,k)}\\ e\sim V^{(2)}\setminus S\end{subarray}}[E_{\delta}(G,S\cup\{e\})^{c}]$
[…]
Expanding the left hand side of Equation 4 and invoking the independence of $G$ and $S$, we obtain failure probability $\varepsilon(-\rho\beta(n)n)$ for $S\sim V^{(2,\varepsilon)}$. ∎
@@ 3.3 Planted Structure
Note that in the proof of Corollary 3.8, we make use the adaptive constants in Theorem 3.6. We can again use this adaptivity to recover variations on existing results in the literature on property resilience (sudakov2008local; frieze2015introduction; bollobas2011random).
@@ 3.4 Reconstruction Consequences
Through Lemma 3.1, a number of reconstruction results follow immediately from Theorems 1.1, 1.2 and 1.3. Using techniques from bollobas1990almost, one can even show that semi-random graphs have reconstruction number three with failure probability $\exp(-\Omega(\beta(n)n))$. Perhaps more interesting is the following characterization of nonreconstructible graphs, which follows from Corollaries 3.10 and 3.11.
@@ 4 Outlook
As noted in the overview, when $S\sim\mathcal{U}(V^{(2)})^{\otimes\varepsilon}$, any reconstructibility result holding with probability 1o(1)1-o(1) and $\varepsilon=\Omega(n^{2}\log(n))$ is equivalent to the deterministic conjecture for large $n$. This leaves a logarithmic gap unknown. Similarly, when $S\sim\mathcal{U}(V^{(2,\varepsilon)})$, it remains unclear whether one can handle any $\varepsilon=n^{2}/2-o(n)$. Such questions outline what is possibly the most natural line for further inquiry.
As suggested in Section 1.1, one can view the semi-random extension of the smoothed model as, instead of uniformly widening the hypercube padding on $\textbf{{p}}\in[0,1]^{N}$, allowing a random (or adversarially chosen) subset $S$ of the entries in this tuple to be arbitrary. One might consider a relaxed setting, where instead of letting $\{p_{i}\}_{i\in S}$ be arbitrary, widening the bound on these entries beyond the $\Theta(\log(n)/n)$ barrier. One may hope to successfully do this on $n^{2}/2-o(n)$ of the entries.
Beyond the semi-random model, there exist several natural notions of approximate graph reconstruction. For instance, take $\operatorname{d\mathfrak{R}}(G)$ to be the minimum number of edge additions/removals necessary to make $G$ reconstructible. One can always isolate a vertex to make $G$ reconstructible, giving $\operatorname{d\mathfrak{R}}(G)\leq n-1$. To the best of our knowledge, however, no sublinear bounds have been established. Interestingly, by strategically changing the degree profile and using myrvold1987bidegreed, one can show that only three edits suffice to make any graph on at least four vertices edge-reconstructible—reconstructible from its multiset of subgraphs with a single edge removed (bondy1991manual). We leave examination of these bounds to future work.
Appendix A Probability
@@ A.1 Negative Association
We closely follow a standard proof for balls and bins. For $e\in V^{(2)}$ and $j\in[\varepsilon]$, let $X_{e,j}$ be a random variable taking the value 1 if $S_{j}=e$ and 0 otherwise. By the Zero-One Lemma (dubhashi1998balls), $\{X_{e,j}\}_{e}$ is NA. Since unions of mutually independent NA families are NA, $\{X_{e,j}\}_{e,j}$ is NA (joag-dev1983negative). But note that $\{\mathbb{I}[e\in S]\}_{e\in V^{(2)}}$ are all monotone increasing in $\{X_{e,j}\}_{e,j}$ and defined on disjoint subsets of $\{X_{e,j}\}_{e,j}$, hence NA. Using the marginal probability bound for NA families,
@@
For $e\in V^{(2)}$, also define $Y_{e}=\mathbb{I}[e\in S]$. Note that $\{X_{j}\}_{j}$ are mutually independent, and that $\{Y_{e}\}_{e\in A\cup B}$ is NA by Lemma A.3. Since these collections are independent, $\mathcal{A}=\{X_{j}\}_{j}\cup\{Y_{e}\}_{e\in A\cup B}$ is NA (joag-dev1983negative). Now observe that
@@ Appendix B Omitted Proofs
$\operatorname*{\mathbb{P}}[\varphi]$ $\leq\exp\left[-\frac{c_{1}}{4}\sum_{i=2}^{N}iM_{i}\right]\leq\begin{cases}\exp\left[-\frac{c_{1}}{2}(n-1)\right]&m=1\\ \exp\left[-\frac{c_{1}}{4}m\left(n-\frac{m}{2}-1\right)\right]&m\geq 2.\end{cases}$
@@
$\leq 4\delta n^{\delta+1}\exp\left[-\frac{c_{1}c_{2}}{2}n\right]$ (5)
@@
$\leq\sum_{m=\lfloor n^{1/2}\rfloor}^{n}\exp\left[m\left(-\frac{\rho^{\prime\prime}}{2}n\right)\right]\leq\exp(\log(n)-\rho^{\prime\prime}n)\leq\exp(-\rho^{\prime}n).$
@@
Appendix C Connection to Deterministic Reconstruction
For completeness, we record connections between deterministic and probabilistic reconstruction statements. We do not pin down precise thresholds.
Observation C.1.
Suppose there exists a family of distributions $\{\mathscr{D}_{n}\}_{1}^{\infty}$ over undirected graphs on $n$ vertices such that for $G\sim\mathscr{D}_{n}$, $S\sim(V^{(2)})^{\otimes\varepsilon}$, and $\varepsilon(n)\geq Cn^{2}\log n$ for $C>1$,
$\operatorname*{\mathbb{P}}[B_{S}(G)\subset\mathfrak{R}_{n}]=1-o(1).$
Then the reconstruction conjecture is true for sufficiently large $n$.
Proof.
Let $H$ be a graph on $n$ vertices. Hence, for $G\sim\mathscr{D}_{n}$, if $H\notin\mathfrak{R}$, then
$\operatorname*{\mathbb{P}}[B_{S}(G)\not\subset\mathfrak{R}\,|\,S=V^{(2)}]=1.$
Expanding, we have
$\operatorname*{\mathbb{P}}[B_{S}(G)\not\subset\mathfrak{R}_{n}]\geq\operatorname*{\mathbb{P}}[S=V^{(2)}]\geq 1-N^{1-C}\geq 1-o(1),$
yielding a contradiction for sufficiently large $n$. ∎
Observation C.2.
Let $\beta\in O(1/n^{2})$ and $\gamma\in o(1)$. Suppose $G\sim\mathscr{G}(n,\textbf{{p}})$ is reconstructible with probability at least $1-\gamma(n)$ for all
$\textbf{{p}}_{i}\in[\beta(n),1-\beta(n)]\quad i\in\binom{n}{2}$.
Then the reconstruction conjecture is true for sufficiently large $n$.
Proof.
Let $H$ be a graph $n$ vertices, and define
$\textbf{{p}}_{i}:=\beta(n)\cdot\mathbb{I}[i\notin E(H)]+(1-\beta(n))\cdot\mathbb{I}[i\in E(H)].$
If $H\notin\mathfrak{R}$, then
$\operatorname*{\mathbb{P}}[G\notin\mathfrak{R}]\geq\operatorname*{\mathbb{P}}[G=H]=\Omega(1).$
Since the left-hand side is $o(1)$, we have a contradiction for sufficiently large $n$. ∎